# The Ginsparg-Wilson relation and overlap fermions

### v1

@@ I Introduction and motivation
The simplest lattice fermions have issues with chiral symmetry. The situation is encoded in a famous “no-go” theorem due to Nielsen and Ninomiya [Nielsen:1980rz, Nielsen:1981xu]. The theorem makes two assumptions:
@@
However there is a third way to evade the theorem, which basically amounts to making a redefinition of what is called “chiral symmetry” on the lattice. Fermion actions which do this are called “Ginsparg-Wilson” actions (after their discoverers [Ginsparg:1981bj]) or “overlap” actions (for their particular implementation choice). They are related to “domain wall fermions,” which live in five dimensions. Domain wall fermions are described by Blum and Shamir in a separate review. The connection is also described in Sec. III of this chapter. Oddly enough, though, nearly all the technical developments of Ginsparg-Wilson fermions and domain wall fermions proceeded in an essentially decoupled way. This chapter attempts to give a stand-alone survey of lattice implementations of overlap fermions, from a description of their formal background to a discussion of issues associated with their practical implementation in simulations.
@@ II Formal properties of Ginsparg-Wilson fermions
Overlap (or, more generally, Ginsparg-Wilson) fermions replace the usual definition of a chiral rotation $\delta\psi=i\epsilon\gamma_{5}\psi$, $\delta\bar{\psi}=i\epsilon\bar{\psi}\gamma_{5}$ by either
@@
This expression is called the “Ginsparg-Wilson relation,” named after its discoverers [Ginsparg:1981bj].
@@
This is the first part of the index theorem, which equates the topological charge (winding number) to $Q$, the difference between zero eigenmodes of positive and negative chirality, gauge configuration by configuration. To complete the theorem it is necessary to relate $Q$ to the topological charge. For a smooth background gauge configuration, showing that $\mbox{Tr}\;\gamma_{5}D$ is proportional to the integral over $\epsilon^{\alpha\beta\mu\nu}F^{a}_{\alpha\beta}F^{a}_{\mu\nu}$ involves manipulations similar to what is done in the continuum calculation (and for that, see the text [Peskin:1995ev]).
@@
With this definition,
@@
Deriving the PCAC relation and Ward identities involves treating the vector and axial currents carefully (see for example, the discussion in Ref. [Hasenfratz:1998jp]), but we can give a “poetic” description of how continuum chiral symmetry is actually obtained. This is due to Niedermeyer [Niedermayer:1998bi]:
In lattice Ward identities the term which breaks the symmetries is the expectation value $\left\langle{\dots\bar{\psi}D\gamma_{5}D\psi\dots}\right\rangle$. The dots label the rest of the expression. The fields $\psi$ and $\bar{\psi}$ then contract against fields in the dots to give $D^{-1}$’s, which in turn contract with the $D$’s in $\left\langle{\dots\bar{\psi}D\gamma_{5}D\psi\dots}\right\rangle$ to give $\left\langle{\dots D^{-1}D\gamma_{5}DD^{-1}\dots}\right\rangle$. This is a contact term (the dots on the left and right sides are pulled to the same point). This means that for all practical purposes, chiral symmetry is exact at nonzero lattice spacing.
Exact chiral symmetry means that the fermion mass is only multiplicatively renormalized by interactions. Another consequence is that lattice to continuum matching factors for suitably chosen vector and axial currents are equal, as are those for the scalar and pseudoscalar currents, $Z_{V}=Z_{A}$, $Z_{P}=Z_{S}$. It’s also the case that $Z_{P}=Z_{S}=1/Z_{m}$ where $Z_{m}$ is the multiplicative matching factor for the fermion mass. (The words “suitably chosen” include the case where the vector current is $V_{\mu}=\bar{q}(x,t)\gamma_{\mu}q(x,t)$, the axial current is $A_{\mu}=\bar{q}(x,t)\gamma_{\mu}\gamma_{5}q(x,t)$, and the scalar and pseudoscalar currents are $S=\bar{q}(x,t)q(x,t)$ and $P=\bar{q}(x,t)\gamma_{5}q(x,t)$, or any other set where the opposite parity terms simply differ by a $\gamma_{5}$.) The PCAC relation $\partial_{\mu}A_{\mu}=2m_{q}P$ plus the definition of the matrix element $\left\langle{0|A_{0}|\pi}\right\rangle=m_{\pi}f_{\pi}$ give $f_{\pi}=(2m_{q}/m_{\pi})\left\langle{0|P|\pi}\right\rangle$ without an additional Z-factor. This is similar to the situation encountered with staggered fermions.
@@
where $V$ is a unitary operator. We can choose $V$ to be a function of some ordinary, nonchiral, undoubled massless lattice Dirac operator $d$ (referred to hereafter as the “kernel” Dirac operator) and a mass term equal to $r_{0}/a$. The universally-used choice of function is the overlap action of Refs. [Neuberger:1997fp, Neuberger:1998my]
@@
All magical things have a high price. Here, the problem is that $V$ is not restricted to extend over a finite number of lattice sites. Said another way, it is not ultralocal. This means that the matrix representation of $D$ is not sparse. It is local: the size of V(|xy|)V(|x-y|) dies exponentially with the separation |xy||x-y|, $V\propto\exp(-C|x-y|)$. Here |xy||x-y| is a distance measured in lattice units; the physical distance is r=a|xy|r=a|x-y| and $V\propto\exp(-Cr/a)$ vanishes when $r\gg a$. The value of $C$ depends on $r_{0}$ and the choice of parameterization for $d$.
@@ III The overlap action beginning in five dimensions
Since the situation does not depend on a particular form for $M$, we will replace it by a wall with Dirichlet boundary conditions at $s=0$. We restrict ourselves to $s>0$ and take $M(s)=M$ in that domain. In this simplified case there is a left-handed zero mode pinned to the $s=0$ surface and its wave function falls away like $\exp(-Ms)$. Now we find the four dimensional effective field theory of the massless mode. We can do this by finding the propagator $G(x,s_{1};y,s_{2})$ between two sites in the five-dimensional manifold and evaluate our result at $s_{1}=s_{2}=0$ (since that is where the massless mode is confined). The Green’s function is the solution to the equation
$D_{5}G(x,s_{1};y,s_{2})=\delta^{4}(x-y)\delta(s_{1}-s_{2})$ (22)
and we impose the boundary condition
$P_{+}G(x,s_{1};y;s_{2})=0$ (23)
at $s_{1}=0$. One can show that the fermion propagator between points on the $s=0$ surface is
$G(x,s_{1}=0;y,s_{2}=0)=2MP_{-}S(x,y)P_{+},$ (24)
where $S(x,y)$ is the inverse of a Dirac operator $D$ given by
$D=M[1-\gamma_{5}\frac{Q}{\sqrt{Q^{2}}}],$ (25)
and where
$Q=\gamma_{5}(M-D_{4}).$ (26)
@@
This, of course is the overlap Dirac operator, with the substitution $M=r_{0}/a$. The derivation of this result was first given by Lüscher [Luscher:2000hn].
To show this, we need the Green’s function ${\cal G}(s_{1},s_{2})$ acting on four-dimensional Dirac fields $\psi(x)$ via
${\cal G}(s_{1},s_{2})=\int d^{4}y\,G(x,s_{1};y,s_{2})\psi(y).$ (27)
$\cal G$} obeys the equation
$D_{5}{\cal G}(s_{1},s_{2})=\delta(s_{1}-s_{2})$ (28)
with the chiral constraint
$P_{+}{\cal G}(s_{1}=0,s_{2})=0.$ (29)
Since $D_{4}$ does not depend on the fifth dimension, we can write down the solution using $Q$ defined above,
${\cal G}(s_{1},s_{2})=e^{(s_{1}-s_{2})Q}[\theta(s_{1}-s_{2})+C]\gamma_{5}$ (30)
before we impose boundary conditions on our solution. Our first constraint is that we want ${\cal G}(s_{1},s_{2})$ to vanish as $|s_{1}-s_{2}|\rightarrow\infty$. We can force that by introducing positive and negative eigenmode projection operators for $Q$, $\hat{P}_{\pm}=\frac{1}{2}(1\pm Q/\sqrt{Q^{2}})\equiv(1\pm\epsilon)/2$, to write
${\cal G}(s_{1},s_{2})=e^{(s_{1}-s_{2})Q}[\theta(s_{1}-s_{2})\hat{P}_{-}-\theta(s_{2}-s_{1})\hat{P}_{+}+C^{\prime}]e^{-s_{2}Q}\gamma_{5}.$ (31)
To make sure that the $C^{\prime}$ term also obeys the boundary conditions, we need to set $C^{\prime}=\hat{P}_{-}O\hat{P}_{+}$ for some $O$. Now we let $s_{1}\rightarrow 0$ and impose the chiral constraint Eq. 29. The identity $2MP_{+}\hat{P}_{-}=P_{+}D$ means that $O=2M/D$. Equation 24 follows from the improbable identities $\epsilon(M/D)=\gamma_{5}(M/D-1)$ and $(M/D)\epsilon\gamma_{5}=\epsilon\gamma_{5}+M/D$. Note that $\hat{P}_{+}+\hat{P}_{-}$ is a contact term, which we will neglect.

### v2

@@ I Introduction and motivation
The simplest lattice fermions have issues with chiral symmetry. The situation is encoded in a famous “no-go” theorem due to Nielsen and Ninomiya [26, 27], see also Karsten and Smit [17]. The theorem makes two assumptions:
@@
However there is a third way to evade the theorem, which basically amounts to making a redefinition of what is called “chiral symmetry” on the lattice. Fermion actions which do this are called “Ginsparg-Wilson” actions (after their discoverers [12]) or “overlap” actions (for their particular implementation choice). They are related to “domain wall fermions,” which live in five dimensions. Domain wall fermions are described by Blum and Shamir in a separate review. The connection is also described in Sec. III of this chapter. Oddly enough, though, nearly all the technical developments of Ginsparg-Wilson fermions and domain wall fermions proceeded in an essentially decoupled way. This chapter attempts to give a stand-alone survey of lattice implementations of overlap fermions, from a description of their formal background to a discussion of issues associated with their practical implementation in simulations.
@@ II Formal properties of Ginsparg-Wilson fermions
Overlap (or, more generally, Ginsparg-Wilson) fermions replace [19] the usual definition of a chiral rotation $\delta\psi=i\epsilon\gamma_{5}\psi$, $\delta\bar{\psi}=i\epsilon\bar{\psi}\gamma_{5}$ by either
@@
This expression is called the “Ginsparg-Wilson relation,” named after its discoverers [12].
@@
This is the first part of the index theorem, which equates the topological charge (winding number) to $Q$, the difference between the numbers of zero eigenmodes of positive and negative chirality, gauge configuration by configuration. To complete the theorem it is necessary to relate $Q$ to the topological charge. For a smooth background gauge configuration, showing that $\mbox{Tr}\;\gamma_{5}D$ is proportional to the integral over $\epsilon^{\alpha\beta\mu\nu}F^{a}_{\alpha\beta}F^{a}_{\mu\nu}$ involves manipulations similar to what is done in the continuum calculation (and for that, see the text [29]).
@@
With this definition, (naturally $H(m)=\gamma_{5}D(m)$),
@@
For the domain wall story about $\hat{D}(m)^{-1}$, see the chapter by Blum and Shamir, Sec. 4.5.
Eq. 17 immediately gives us $m_{PS}^{2}\propto m_{q}$: multiply both sides on the left by $\gamma_{5}$, trace and sum over sites,
$2{\rm Tr}\hat{D}^{-1}=2m{\rm Tr}\gamma_{5}\hat{D}(m)^{-1}\gamma_{5}\hat{D}(m)^{-1}.$ (18)
The right hand side has the topology of a flavor nonsinglet pseudoscalar correlator. Eq. 18 can be written as
$\frac{1}{V}\sum_{x}\left\langle{S(x)}\right\rangle=\frac{m_{q}}{V}\sum_{x,y}\left\langle{P(x)P(y)}\right\rangle$ (19)
with the scalar and pseudoscalar currents defined (to leading order in $m$) as
$S=\bar{q}(x,t)(1-\frac{a}{2r_{0}}D)q(x,t);\quad P=\bar{q}(x,t)(1-\frac{a}{2r_{0}}D)\gamma_{5}q(x,t).$ (20)
When chiral symmetry is broken, the left hand side of Eq. 19 is the chiral condensate $\Sigma$. Saturating the correlator on the right hand side of Eq. 19 with the pseudoscalar state gives us
$C(t)=\sum_{\vec{x}}\left\langle{P(\vec{x},t)P(0,0)}\right\rangle=|\left\langle{0|P|PS}\right\rangle|^{2}\frac{\cosh(m_{PS}(N_{t}/2-t))}{2m_{PS}\sinh(m_{PS}N_{t}/2)}.$ (21)
and its integral over the simulation volume is
$\int_{0}^{N_{t}}dtC(t)=\frac{|\left\langle{0|P|PS}\right\rangle|^{2}}{m_{PS}^{2}}.$ (22)
When $\Sigma$ is nonzero this gives us
$\frac{|\left\langle{0|P|PS}\right\rangle|^{2}}{m_{PS}^{2}}=\frac{\Sigma}{m_{q}}.$ (23)
which is $m_{PS}^{2}\propto m_{q}$.
Chiral Ward identities come from writing Eq. 17 as
$\hat{D}(m)^{-1}+\gamma_{5}\hat{D}(m)^{-1}\gamma_{5}=2m\gamma_{5}\hat{D}(m)^{-1}\gamma_{5}\hat{D}(m)^{-1}$ (24)
so
${\rm Tr}\hat{D}(m)^{-1}\Gamma\hat{D}(m)^{-1}\Gamma+{\rm Tr}\hat{D}(m)^{-1}\Gamma(\gamma_{5}\hat{D}(m)^{-1}\gamma_{5})\Gamma$
$=2m{\rm Tr}\hat{D}(m)^{-1}\Gamma(\gamma_{5}\hat{D}(m)^{-1}\gamma_{5}\hat{D}(m)^{-1}\gamma_{5})\Gamma$
for any gamma matrix $\Gamma$. For $\Gamma=\gamma_{\mu}$ this is the vector - axial vector - pseudoscalar Ward identity
$\left\langle{V_{\mu}V_{\mu}}\right\rangle-\left\langle{A_{\mu}A_{\mu}}\right\rangle=2m\left\langle{A_{\mu}PV_{\mu}}\right\rangle.$ (26)
for
$V_{\mu}=\bar{q}(x,t)(1-\frac{a}{2r_{0}}D)\gamma_{\mu}q(x,t);\quad A_{\mu}=\bar{q}(x,t)(1-\frac{a}{2r_{0}}D)\gamma_{\mu}\gamma_{5}q(x,t).$ (27)
None of the currents we have defined are Noether currents, so there will be matching factors connecting them to their continuum analogs. The continuum Ward identities are identical to the ones we have written. Eq. 26, plus the corresponding one with $\Gamma=1$, tell us that lattice to continuum matching factors for the vector and axial currents are equal, as are those for the scalar and pseudoscalar currents, $Z_{V}=Z_{A}$, $Z_{P}=Z_{S}$.
Notice that the extra terms in the currents and identities compared to the continuum relations involve the product $D^{-1}D$, which is a contact term in the correlation functions.
The singlet analog of Eq. 18 includes the topological charge and gives the Di Vecchia and Veneziano [8], Leutwyler and Smilga [18], and Crewther [4] relation, that the topological susceptibility vanishes linearly with the fermion mass. For details, see (for example) the early review by Niedermeyer [25].
@@
where $V$ is a unitary operator. We can choose $V$ to be a function of some ordinary, nonchiral, undoubled massless lattice Dirac operator $d$ (referred to hereafter as the “kernel” Dirac operator) and a mass term equal to $r_{0}/a$. The universally-used choice of function is the overlap action of Refs. [24, 23]
@@
All magical things have a high price. Here, the problem is that $V$ is not restricted to extend over a finite number of lattice sites. Said another way, it is not ultralocal. This means that the matrix representation of $D$ is not sparse.
There was a theoretical issue in the literature: was $D$ local? “Locality” means that the size of D(|xy|)D(|x-y|) dies exponentially with the separation |xy||x-y|, $D\propto\exp(-C|x-y|)$. Here |xy||x-y| is a distance measured in lattice units; the physical distance is r=a|xy|r=a|x-y| and $D\propto\exp(-Cr/a)$ vanishes when $r\gg a$. The answer to this question depended on the choice of the kernel. A proof of locality when the kernel $d$ was the Wilson action with thin links, but in the limit of very smooth gauge fields, was presented in Ref. [14]. Most of the literature involved issues related to using the Wilson action as a kernel. I never did that, so for me the only issue of principle is illustrated in Fig. 1. The overlap maps $d$ into a circle. All of the eigenmodes of $d$ with real eigenvalues closer to the origin than the value of $r_{0}/a$ will be pushed to the origin and act as separate flavors. We don’t want doubling, so there had better only be one set of these modes. In the figure, it had better be that $r_{0}/a<1$. This was never an issue in practice to maintain. Also, in practice, $C$ was never known from first principles, so everybody who ever simulated overlap fermions looked at a plot of something like $D(x-y)\delta(y)$ as a function of |xy||x-y| and tuned $d$ and $r_{0}$ to minimize the range of $D$.
@@ III The overlap action beginning in five dimensions
Since the situation does not depend on a particular form for $M$, let us replace it by a wall with Dirichlet boundary conditions at $s=0$. We restrict ourselves to $s>0$ and take $M(s)=M$ in that domain. For this simplified case the left-handed zero mode is pinned to the $s=0$ surface and its wave function falls away like $\exp(-Ms)$. Now let us find the four dimensional effective field theory of the massless mode. We can do this by finding the propagator $G(x,s;y,s_{0})$ between two sites in the five-dimensional manifold and restrict our attention to the case $s=s_{0}=0$ (since that is where the massless mode is confined).
The Green’s function obeys
$D_{5}G(x,s;y,s_{0})=\delta^{4}(x-y)\delta(s-s_{0})$ (32)
with the boundary conditions that $G(x,s;y,s_{0})$ should vanish as $|s-s_{0}|$ goes to infinity and
$P_{+}G(x,s;y;s_{0})=0$ (33)
at $s=0$. Here I introduce the chiral projectors
$P_{\pm}=\frac{1}{2}(1\pm\gamma_{5}).$ (34)
A few more definitions are useful before we proceed. First, set
$Q=\gamma_{5}(M-D_{4})$ (35)
so
$D_{5}=\gamma_{5}(\partial_{5}-Q).$ (36)
To make contact with the notation of previous sections, $M=r_{0}/a$, $D_{4}=d$ is the kernel operator and $Q=-h(-r_{0}/a)$ is the Hermitian Dirac operator. Eigenmodes of $Q$,
$Q\mathinner{|{n}\rangle}=q_{n}\mathinner{|{n}\rangle},$ (37)
will be used to construct the solution. Positive and negative eigenvalue projectors of $Q$,
$\hat{P}_{\pm}$ $=$ $\frac{1}{2}(1\pm\frac{Q}{\sqrt{Q^{2}}})$
$\equiv$ $\frac{1}{2}(1\pm\epsilon(Q)),$
will play a rôle in what follows.
Our goal is to show that
$G(x,s=0;y,s_{0}=0)=2MP_{-}S(x,y)P_{+},$ (39)
where $S(x,y)=D_{ov}(x,y)^{-1}$ if $x\neq y$, with
$D_{ov}=M[1-\gamma_{5}\epsilon(Q)].$ (40)
@@
That is, $D_{ov}$ is the overlap Dirac operator. The derivation of this result was first given by Lüscher [20].
To begin, write the four dimensional delta function in terms of the eigenmodes of $Q$
$\delta^{4}(x-y)=\sum_{n}\left\langle{x|n}\right\rangle\left\langle{n|y}\right\rangle$ (41)
and likewise write the Green’s function as
$G(s,x;s_{0},y)=\sum_{n}\left\langle{x|n}\right\rangle{\cal G}_{n}(s,s_{0})\left\langle{n|y}\right\rangle.$ (42)
Then ${\cal G}_{n}(s,s_{0})$ obeys the differential equation
$\gamma_{5}(\partial_{5}-q_{n}){\cal G}_{n}(s,s_{0})=\delta(s-s_{0}).$ (43)
Leaving aside the boundary conditions for the moment, the solution to this simple problem is
${\cal G}_{n}(s,s_{0})$ $=$ $[Ae^{q_{n}(s-s_{0})}]\gamma_{5}\qquad\rm{if}\ s<s_{0}$
$=$ $[Be^{q_{n}(s-s_{0})}]\gamma_{5}\qquad\rm{if}\ s>s_{0}$
where BA=1B-A=1 to handle the discontinuity due to the delta function. At this point we can write the Green’s function as
$G(s,x;s_{0},y)=\sum_{n}\left\langle{x|n}\right\rangle e^{q_{n}s}[\theta(s-s_{0})+C]e^{-q_{n}s_{0}}\gamma_{5}\left\langle{n|y}\right\rangle.$ (45)
(I have split the $\left\langle{x|n}\right\rangle$ and $\left\langle{n|y}\right\rangle$ apart because at the end I will want to write the answer directly in terms of operators depending on $Q$.)
Now for the boundary conditions. We want the solution to fall to zero as $|s-s_{0}|$ goes to infinity. Make the replacement
[…]
in Eq. 45. The $\hat{P}_{-}$ part of this expression gives a contribution to the Green’s function which falls to zero at big $|s-s_{0}|$. To handle the positive eigenmodes, write
$\theta(s-s_{0})\hat{P}_{+}=-(1-\theta(s-s_{0}))\hat{P}_{+}+\hat{P}_{+}$. (47)
Now we have
$G(s,x;s_{0},y)$ $=$ $\sum_{n}\left\langle{x|n}\right\rangle e^{q_{n}s}[\theta(s-s_{0})\hat{P}_{-}$
$-(1-\theta(s-s_{0}))\hat{P}_{+}+(\hat{P}_{+}+C)]e^{-q_{n}s_{0}}\gamma_{5}\left\langle{n|y}\right\rangle.$
or
$G(s,x;s_{0},y)=\sum_{n}\left\langle{x|n}\right\rangle e^{q_{n}s}[\theta(s-s_{0})\hat{P}_{-}-\theta(s_{0}-s)\hat{P}_{+}+C^{\prime}]e^{-q_{n}s_{0}}\gamma_{5}\left\langle{n|y}\right\rangle.$ (49)
The first two terms satisfy the desired boundary condition at large $|s-s_{0}|$. To make sure that the $C^{\prime}$ term also obeys the boundary condition, we need $C^{\prime}=\hat{P}_{-}O\hat{P}_{+}$ for some $O$. We have arrived at
$G(s,x;s_{0},y)$ $=$ $\sum_{n}\left\langle{x|n}\right\rangle e^{q_{n}s}[\theta(s-s_{0})\hat{P}_{-}$
$-\theta(s_{0}-s)\hat{P}_{+}+\hat{P}_{-}O\hat{P}_{+}]e^{-q_{n}s_{0}}\gamma_{5}\left\langle{n|y}\right\rangle.$
Now we take the limit $s\rightarrow 0$ first, then $s_{0}\rightarrow 0$. In everything we have done so far, the $\hat{P}_{\pm}$’s were functions of the $q_{n}$’s, but we can now interpret them as functions of the $Q$’s (which are themselves functions of the four dimensional coordinates) and strip off the explicit dependence on the eigenmodes of $Q$. Suppressing the labels $s$ and $s_{0}$, our Green’s function is
$G(x,y)=[-\hat{P}_{+}+\hat{P}_{-}O\hat{P}_{+}]\gamma_{5}.$ (51)
On to the last boundary condition, Eq. 33. This is
$(-P_{+}+P_{+}\hat{P}_{-}O)\hat{P}_{+}=0$ (52)
We use
$P_{+}\hat{P}_{-}=\frac{1}{2}(1+\gamma_{5})\frac{1}{2}(1-\epsilon)=\frac{1}{2}(1+\gamma_{5})\frac{1}{2}(1-\gamma_{5}\epsilon)=P_{+}\frac{D_{ov}}{2M}$ (53)
which says that
$O=\frac{2M}{D_{ov}}.$ (54)
The Green’s function is
$G(x,y)=[-\hat{P}_{+}+\hat{P}_{-}\frac{2M}{D_{ov}}\hat{P}_{+}]\gamma_{5}.$ (55)
Finally, the improbable identities (for the student to check!) $\epsilon(Q)(M/D_{ov})=\gamma_{5}(M/D_{ov}-1)$, $(M/D_{ov})\epsilon(Q)\gamma_{5}=\epsilon\gamma_{5}+M/D_{ov}$, and $\epsilon(Q)(M/D_{ov})\epsilon(Q)\gamma_{5}=\gamma_{5}(M/D_{ov})$ give
$G(x,y)=P_{-}+P_{-}\frac{2M}{D_{ov}}P_{+}.$ (56)
which is what we wanted to show. The $P_{-}$ term is a contact term, and we can neglect it.